Find all real and imaginary solutions to the equation.
b^4+13b^2+36=0

Answer by checkley75(3666)

B^4+13B^2+36=0
(B^2+9)(B^2+4)=0
B^2=-9
B=+-3i ANSWER.
B^2+4=0
B^2=-4
B=+-2i ANSWER